I was 28 years old when my son was born. Today, two times the difference of my age and four equals 10 times my son's age now. How old are we today?
Answer by stanbon(75887)

I was 28 years old when my son was born. Today, two times the difference of my age and four equals 10 times my son's age now. How old are we today?
----------
Equations:
f = s + 28
2(f-4)= 10s
---------
Modify and substitute for "f":
2(s+28 -4) = 10s
2s + 48 = 10s
8s = 48
son's age now = 6 yrs
----
f = s + 28
f = 34
father's age now is 34
----------------------------
Cheers,
Stan H.
=================